His4life:

Equal weights of Zn metal and iodine are mixed together and is completely converted to. What fraction by weight of original Zn remains unreacted? (Zn = 65, I = 127)

toga:

The balanced chemical equation for the reaction between Zn and I2 is: Zn + I2 → ZnI2 The molar mass of Zn is 65 g/mol and the molar mass of I2 is 254 g/mol. If equal weights of Zn and I2 are mixed together, then the number of moles of each reactant is: Number of moles of Zn = Mass of Zn / Molar mass of Zn = (65 g) / (65 g/mol) = 1 mol Number of moles of I2 = Mass of I2 / Molar mass of I2 = (65 g) / (254 g/mol) = 0.256 mol According to the balanced chemical equation, 1 mol of Zn reacts with 1 mol of I2 to form 1 mol of ZnI2. Therefore, the number of moles of Zn that react is 0.256 mol, and the number of moles of Zn that remain unreacted is 1 - 0.256 = 0.744 mol. The mass of the unreacted Zn is: Mass of unreacted Zn = Number of moles of unreacted Zn x Molar mass of Zn = (0.744 mol) x (65 g/mol) = 48.36 g Therefore, the fraction by weight of original Zn that remains unreacted is: Fraction by weight of unreacted Zn = Mass of unreacted Zn / Total mass of Zn and I2 = (48.36 g) / (65 g + 65 g) = 0.372 or 37.2%

His4life:

Wasn't expecting that much but ok the answer I got was 0.74=74%

toga:

@his4life wrote:
Wasn't expecting that much but ok the answer I got was 0.74=74%
I like to show my work

His4life:

@toga wrote:
@his4life wrote:
Wasn't expecting that much but ok the answer I got was 0.74=74%
I like to show my work
oh ok

Luigi0210:

@his4life wrote:
Wasn't expecting that much but ok the answer I got was 0.74=74%
You would be right Is this what you ended up with? 1 mol of \(\large I_2\) and 3.88 mol of Zn So 2.88 mol of excess Zn, equating to: \(\Large \frac{(65.39*2.88)}{254} =0.7370 *100\approx \) 74%

His4life:

Yeah that's what I ended up with

Luigi0210:

Then yes, seems right. The way Toga did it, there would be no excess Zn, as it would all be used up

His4life:

oh ok